# Foundation Model for Polycrystalline Material Informatics

### v2

@@ Abstract
We present a three-dimensional polycrystal foundation model based on a masked autoencoder that learns intrinsic microstructural representations through large-scale self-supervised pretraining on voxel-based data. The pretraining dataset consists of 100,000100{,}000 face-centered cubic (FCC) microstructures whose crystallographic textures span the texture hull via hierarchical simplex sampling. The quality and transferability of the learned representations are evaluated through two downstream tasks: (i) homogenized stiffness prediction and (ii) nonlinear homogenized response prediction. In the latter, the pretrained encoder is coupled with an orientation-aware interaction-based deep material network (ODMN), where the learned latent representations are used to infer microstructure-dependent ODMN parameters. This enables accurate stress–strain predictions for previously unseen microstructures under crystal plasticity. Across both tasks, the pretrained encoder consistently exhibits superior generalization performance compared to non-pretrained baselines. These results demonstrate the strong transferability of the proposed foundation model and its effectiveness in data-scarce scientific settings with limited labeled microstructures. The framework further enables scalable integration with experimentally derived microstructures, providing a practical basis for microstructure–property reasoning in materials design.
@@ 1 Introduction
Polycrystalline materials form the backbone of modern engineering applications, with their macroscopic mechanical behavior intrinsically governed by the underlying microstructural architecture, most notably the crystallographic texture [motaman2020anisotropic]. As a result, elucidating and quantifying microstructure–property relationships has become a central objective in materials design and performance optimization. In recent years, the rapid advancement of artificial intelligence (AI) techniques has opened new avenues for establishing data-driven links between microstructural descriptors and effective material responses, offering a promising route toward accelerated computational materials modeling.
@@
Within this context, a wide range of machine learning architectures have been explored for polycrystalline systems. Convolutional neural networks (CNNs), such as U-Net, have been employed to predict full-field stress distributions in viscoplastic polycrystals [khorrami2023artificial]. Graph neural networks (GNNs) have demonstrated strong capability in capturing grain-level topological interactions, enabling accurate prediction of effective magnetostriction in heterogeneous polycrystals [dai2021graph]. Variational autoencoders (VAEs) have been used to learn compact latent representations of electron backscatter diffraction (EBSD) patterns, yielding physically meaningful embeddings that improve EBSD indexing efficiency [liu2025learning]. More recently, masked autoencoders (MAEs) have been applied to synthetic two-dimensional polycrystalline datasets for classification tasks; although the pretraining is restricted to random textured microstructures, the results nevertheless highlight the potential of self-supervised learning for texture-informed representation learning [belamri2025quaternion]. Collectively, these studies underscore the critical role of latent representation learning in advancing microstructure-aware materials modeling.
@@
Building upon this progress, the concept of a foundation model extends representation learning by leveraging large-scale pretraining on abundant data to extract universal latent representations that are transferable across diverse downstream tasks. In natural language processing, large language models such as BERT and GPT exemplify this paradigm by learning highly generalizable embeddings from massive text corpora [devlin2018bert, radford2019language]. Analogously, foundation models tailored to materials science have begun to emerge. For example, a masked autoencoder pretrained on synthetic two-dimensional composite microstructures has demonstrated strong versatility across downstream tasks, including nonlinear stress–strain response prediction [WEI2025114397].
@@
Motivated by these developments, this study explores the potential of foundation models as a unifying framework for crystallographic texture informatics in polycrystalline materials. To this end, we pretrain a foundation model on a large-scale synthetic dataset that systematically spans the crystallographic texture space. The dataset is constructed using hierarchical simplex sampling (HSS) [johnson2018efficient, nino2024evolution] to traverse the complete texture hull of face-centered cubic (FCC) crystals, thereby capturing the full diversity of crystallographic texture states. Voxel-based polycrystalline microstructures are generated from these sampled textures and used for large-scale self-supervised pretraining.
@@
Building upon this dataset, we develop a three-dimensional voxel-based foundation model for polycrystalline material informatics based on a masked autoencoder architecture, which learns texture-aware latent representations in a self-supervised manner. The pretrained model is subsequently fine-tuned for two representative downstream tasks: (1) prediction of homogenized stiffness, and (2) inference of material surrogate model parameters via a linear projection of the learned latent representations. In the second task, the inferred surrogate parameters are evaluated under crystal plasticity to predict nonlinear homogenized responses. Across both tasks, the pretrained model consistently outperforms non-pretrained baselines, demonstrating strong generalization capability and underscoring the effectiveness of large-scale self-supervised pretraining for texture-aware representation learning.
@@
The remainder of this paper is organized as follows. Section 2 details the proposed methodology, including synthetic dataset generation, foundation model pretraining, and downstream fine-tuning strategies. Section 3 presents and discusses the numerical results for both homogenized stiffness prediction and nonlinear response modeling. Finally, Section 4 summarizes the key findings and outlines perspectives for future developments in foundation model–based approaches for polycrystalline materials.
@@ 2.2 Self-supervised pretraining stage
To encode spatial information, three-dimensional sinusoidal positional encodings (PEs) are added to the inputs of both the encoder and decoder following the formulations in [vaswani2017attention, WEI2025114397]. After patchification, the patch grid has dimensions $\mathcal{G}\times\mathcal{G}\times\mathcal{G}$, with $\mathcal{G}=5$ in this study. Each spatial coordinate $(x,y,z)$ is independently encoded using sinusoidal functions at multiple frequencies:
$\text{PE}_{x}^{(2i)}(x)$ $=\sin\left(\frac{x}{10000^{2i/\mathcal{D}_{e}}}\right),\quad\text{PE}_{x}^{(2i+1)}(x)=\cos\left(\frac{x}{10000^{2i/\mathcal{D}_{e}}}\right),$ (3)
$\text{PE}_{y}^{(2i)}(y)$ $=\sin\left(\frac{y}{10000^{2i/\mathcal{D}_{e}}}\right),\quad\text{PE}_{y}^{(2i+1)}(y)=\cos\left(\frac{y}{10000^{2i/\mathcal{D}_{e}}}\right),$
$\text{PE}_{z}^{(2i)}(z)$ $=\sin\left(\frac{z}{10000^{2i/\mathcal{D}_{e}}}\right),\quad\text{PE}_{z}^{(2i+1)}(z)=\cos\left(\frac{z}{10000^{2i/\mathcal{D}_{e}}}\right),$
where $i=0,1,\ldots,\mathcal{D}_{e}/2-1$.
@@
$\mathbf{PE}_{\text{encoder}}=\text{stack}\left(\text{PE}_{\text{[CLS]}},\,\left\{\text{concat}\left(\text{PE}_{x}(x),\,\text{PE}_{y}(y),\text{PE}_{z}(z)\right)\mid(x,y,z)\right\}\right)\in\mathbb{R}^{(1+N_{p})\times\mathcal{D}_{e}}$ (4)
@@ 2.3.1 Downstream task I: Homogenized stiffness prediction
2.3.2 Overview of the ODMN
Accurately predicting the nonlinear mechanical response of RVEs remains a long-standing challenge, particularly in industrial applications where direct full-field simulations are computationally prohibitive. To address this challenge, the orientation-aware interaction-based deep material network (ODMN) is proposed as an efficient and physically interpretable surrogate modeling framework [Wei01]. The ODMN constitutes an extension of the deep material network (DMN) architecture [liu2019deep, noels2022micromechanics, noels2022interaction], and is specifically tailored for polycrystalline materials to explicitly incorporate crystallographic texture effects.
Conceptually, the ODMN can be interpreted as a parameterized reduced-order microstructure that approximates the homogenized mechanical response of the original RVE [wu2025stochastic, shin2023deep]. Starting from a unit cell $\Omega$, the reduced-order microstructure is hierarchically partitioned into $2^{N}$ subdomains,
@@ 2.3.2 Overview of the ODMN
where each subdomain $\Omega_{i}$ corresponds to a material node $\mathcal{M}^{i}$ in the network. Here, $N$ is a user-defined hyperparameter that controls the resolution of the reduced-order representation, with larger values of $N$ enabling a richer approximation of the RVE homogenized response.
The hierarchical partitioning is encoded through a binary-tree material network of depth $N$. At each internal tree node located at level $l$ and position $p$, an interaction direction $\vec{\mathbf{N}}^{\,l}_{p}$ is introduced to characterize the interface orientation governing stress equilibrium between the two adjacent subdomains. At each leaf-level material node $\mathcal{M}^{i}$, a set of parameters $\{z^{i},\alpha^{i},\beta^{i},\gamma^{i}\}$ is assigned, representing the volume fraction and the local crystallographic orientation associated with subdomain $\Omega_{i}$.
The interaction direction $\vec{\mathbf{N}}$ is parameterized by two angular variables, $\theta$ and $\phi$, which define a unit normal vector according to
@@
For an ODMN of depth $N$, the complete set of trainable parameters can be expressed as
@@
where the first set corresponds to the material-node parameters associated with the reduced subdomains, and the second set specifies the interaction directions at each tree node of the hierarchical material network.
A distinctive feature of the ODMN, inherited from the DMN framework, is that the network can be trained in an offline stage using only linear elastic stiffness data, while still being capable of performing nonlinear mechanical extrapolation during the online prediction stage [liu2019deep, gajek2020micromechanics, noels2022micromechanics, wan2024decoding].
This remarkable mechanical extrapolation capability originates from the fact that the ODMN parameters learned during offline training encode a first-order approximation of the microstructural mechanical response under linear elasticity. During online prediction, this first-order approximation is preserved and embedded within the hierarchical homogenization operators, thereby enabling physically consistent extrapolation to nonlinear mechanical behavior [gajek2020micromechanics].
During the online prediction stage, the ODMN performs a downscaling operation from the macroscopic deformation gradient $\bar{\mathbf{F}}$ to the deformation gradient $\mathbf{F}^{i}$ associated with each subdomain $\Omega_{i}$ according to
@@
where the interaction coefficients $\alpha^{i,j}$ and the corresponding interaction directions $\mathbf{N}^{j}$ are determined by the network parameter set $\mathcal{F}$.
At each material node, the local constitutive response is evaluated through the material law
@@
where $\mathbf{P}^{i}$ denotes the first Piola–Kirchhoff stress of subdomain $\Omega_{i}$, $\partial\mathbf{P}^{i}/\partial\mathbf{F}^{i}$ is the corresponding consistent tangent operator, and $\mathbb{P}^{i}(\cdot)$ represents the local material law parameterized by the internal state variables $\boldsymbol{\psi}^{i}$ (e.g., history-dependent variables in crystal plasticity).
The resulting local stresses and tangents are subsequently upscaled through the hierarchical network structure to recover the homogenized first Piola–Kirchhoff stress $\bar{\mathbf{P}}$ and its consistent tangent $\partial\bar{\mathbf{P}}/\partial\bar{\mathbf{F}}$.
The interaction vectors $\mathbf{a}^{j}$ introduced in Eq. (9) represent fluctuation fields associated with the interfaces between adjacent subdomains in the reduced-order microstructure. During each Newton–Raphson iteration of the online prediction stage, the interaction vectors are initialized as $\mathbf{a}^{j}=\mathbf{0}$. After evaluating the local constitutive responses at the material nodes, the consistency of the downscaled and upscaled responses is assessed through the Hill–Mandel condition, which requires
@@
where $W^{i}$ denotes the volume fraction associated with subdomain $\Omega_{i}$.
If the Hill–Mandel condition is violated, the interaction vectors $\mathbf{a}^{j}$ are updated via a Newton–Raphson procedure, and the downscaling–local evaluation–upscaling cycle is repeated until convergence is achieved. Upon convergence, the homogenized first Piola–Kirchhoff stress $\bar{\mathbf{P}}$ and its consistent tangent $\partial\bar{\mathbf{P}}/\partial\bar{\mathbf{F}}$ are obtained and used to define the macroscopic constitutive response. Further details can be found in the original ODMN formulation [Wei01].
Similar to the original DMN architecture, the ODMN exhibits an inherent limitation in that its parameters are tied to a specific microstructural realization, thereby restricting transferability across different microstructures. To address this limitation, several recent studies have proposed strategies to generalize DMN parameters through interpolation schemes or machine-learning-based frameworks. For example, Huang et al. [huang2022microstructure] developed an interpolation-based scheme to predict DMN parameters for previously unseen microstructures. In contrast, Jean et al. [jean2024graph] employed graph neural networks (GNNs) to infer DMN parameters directly from mesh-based microstructural representations. Wu et al. [wu2026convolutional] utilized convolutional neural networks (CNNs) to map two-dimensional microstructural images into a latent feature space, followed by a feedforward neural network to predict the corresponding DMN parameters. More recently, Wei et al. [WEI2025114397] demonstrated that foundation models pretrained on large-scale short-fiber composite images can serve as universal feature extractors, whose learned latent representations are subsequently mapped to DMN parameters.
Nevertheless, the extension of such approaches to three-dimensional polycrystalline microstructures, particularly for mapping learned representations to ODMN parameters, remains largely unexplored. Motivated by this gap, the present study investigates how a foundation model can act as a latent feature extractor for polycrystalline microstructures, whose learned representations are subsequently mapped to surrogate models, such as the ODMN, to enable generalization across diverse microstructural configurations.

### v3

@@ Abstract
We present a three-dimensional foundation model for polycrystalline materials based on a masked autoencoder trained via large-scale self-supervised learning. The model is pretrained on 100,000100{,}000 voxelized synthetic face-centered cubic (FCC) microstructures whose crystallographic textures systematically span the texture hull using hierarchical simplex sampling. The transferability of the learned latent representations is evaluated on two downstream tasks: homogenized elastic stiffness prediction and nonlinear stress-strain response prediction. For the nonlinear task, the pretrained encoder is coupled with an orientation-aware interaction-based deep material network (ODMN), where latent features are used to infer microstructure-dependent surrogate parameters. The inferred ODMNs are subsequently combined with crystal plasticity to predict stress–strain responses for previously unseen microstructures. In stiffness prediction, the pretrained model achieves validation $R^{2}$ values exceeding 0.8, compared to below 0.1 for non-pretrained baselines. In nonlinear response prediction, mean stress errors remain below 4%. These results demonstrate that self-supervised pretraining yields physically meaningful and transferable microstructural representations, providing a scalable framework for microstructure-property inference in polycrystalline materials.
@@ 1 Introduction
Polycrystalline materials form the backbone of modern engineering applications, with their macroscopic mechanical behavior intrinsically governed by the underlying microstructural architecture, most notably crystallographic texture [motaman2020anisotropic]. Consequently, elucidating microstructure–property relationships has become a central objective in materials design and performance optimization. In recent years, advances in artificial intelligence (AI) have opened new opportunities for establishing data-driven links between microstructural descriptors and effective material responses, offering a promising route toward accelerated computational materials modeling.
@@
A wide range of machine learning architectures have been explored for polycrystalline systems. Convolutional neural networks (CNNs), such as U-Net, have been employed to predict full-field stress distributions in viscoplastic polycrystals [khorrami2023artificial]. Graph neural networks (GNNs) have demonstrated strong capability in capturing grain-level topological interactions, enabling accurate prediction of effective magnetostriction in heterogeneous polycrystals [dai2021graph]. Variational autoencoders (VAEs) have been used to learn compact latent representations of electron backscatter diffraction (EBSD) patterns, yielding physically meaningful embeddings that improve EBSD indexing efficiency [liu2025learning]. More recently, masked autoencoders (MAEs) have been applied to synthetic two-dimensional polycrystalline datasets for classification tasks; although pretraining was restricted to randomly textured microstructures, these studies nevertheless highlight the potential of self-supervised learning for texture-aware representation learning [belamri2025quaternion]. Collectively, these efforts underscore the importance of latent representation learning in microstructure-aware materials modeling.
@@
Building upon this progress, the concept of a foundation model extends representation learning by leveraging large-scale pretraining on abundant data to extract transferable latent representations across diverse downstream tasks. In natural language processing, large language models such as BERT and GPT exemplify this paradigm by learning highly generalizable embeddings from massive text corpora [devlin2018bert, radford2019language]. Analogously, foundation models tailored to materials science have begun to emerge. For example, a masked autoencoder pretrained on synthetic two-dimensional composite microstructures has demonstrated strong versatility across downstream tasks [WEI2025114397].
To efficiently predict nonlinear mechanical responses of polycrystalline materials, physically informed surrogate models have been developed as alternatives to direct full-field simulations. Among these approaches, the orientation-aware interaction-based deep material network (ODMN) has been proposed as an efficient and physically interpretable surrogate modeling framework [Wei01]. The ODMN extends the deep material network (DMN) architecture [liu2019deep, noels2022micromechanics, noels2022interaction] and explicitly incorporates crystallographic texture effects through a hierarchical reduced-order representation. A distinctive feature of the ODMN is that its parameters can be trained using linear elastic homogenized responses, while remaining capable of extrapolating to nonlinear mechanical behavior during online prediction [liu2019deep, gajek2020micromechanics, noels2022micromechanics, wan2024decoding]. However, ODMN parameters are tied to individual microstructural realizations, limiting transferability across different textures. To address this issue, several studies have proposed interpolation- or learning-based strategies to generalize DMN parameters across microstructures [huang2022microstructure, jean2024graph, wu2026convolutional, WEI2025114397]. Nevertheless, extension of such approaches to three-dimensional polycrystalline microstructures, particularly for mapping learned representations to ODMN parameters, remains largely unexplored.
@@
Motivated by these developments, this study investigates foundation models as a unifying framework for crystallographic texture informatics in polycrystalline materials. We introduce a foundation model as a universal microstructure encoder, whose learned latent representations are systematically mapped to ODMN parameters, enabling transferable nonlinear response prediction for previously unseen microstructures. The model is pretrained on a large-scale synthetic dataset spanning the complete crystallographic texture space of face-centered cubic (FCC) crystals, generated via hierarchical simplex sampling (HSS) [johnson2018efficient, nino2024evolution]. Voxel-based polycrystalline microstructures constructed from these textures are used for large-scale self-supervised pretraining.
@@
Building upon this dataset, we develop a three-dimensional voxel-based foundation model based on a masked autoencoder architecture, which learns texture-aware latent representations in a self-supervised manner. The pretrained model is subsequently fine-tuned for two representative downstream tasks: (1) prediction of homogenized stiffness and (2) inference of material surrogate model parameters via linear projection of the learned latent representations. In the second task, the inferred surrogate parameters are evaluated under crystal plasticity to predict nonlinear homogenized responses. Across both tasks, the pretrained model consistently outperforms non-pretrained baselines, demonstrating improved generalization and highlighting the effectiveness of large-scale self-supervised pretraining for texture-aware representation learning.
@@
The remainder of this paper is organized as follows. Section 2 details the proposed methodology, including dataset generation, foundation model pretraining, and downstream fine-tuning strategies. Section 3 presents and discusses the numerical results for homogenized stiffness prediction and nonlinear response modeling. Finally, Section 4 summarizes the key findings and outlines perspectives for future developments in foundation model-based approaches for polycrystalline materials.
@@ 2.2 Self-supervised pretraining stage
To encode spatial information, three-dimensional sinusoidal positional encodings (PEs) are added to the inputs of both the encoder and decoder following the formulations in [vaswani2017attention, WEI2025114397]. After patchification, the patch grid has dimensions $\mathcal{G}\times\mathcal{G}\times\mathcal{G}$, with $\mathcal{G}=5$ in this study. Each spatial coordinate $(x,y,z)$ is independently encoded using sinusoidal functions at multiple frequencies. The embedding dimension $\mathcal{D}_{e}$ is evenly partitioned across the three spatial axes, such that $\mathcal{D}_{x}=\mathcal{D}_{y}=\mathcal{D}_{z}=\mathcal{D}_{e}/3$. The axis-wise positional encodings are defined as
$\text{PE}_{x}^{(2i)}(x)$ $=\sin\left(\frac{x}{10000^{2i/\mathcal{D}_{x}}}\right),\quad\text{PE}_{x}^{(2i+1)}(x)=\cos\left(\frac{x}{10000^{2i/\mathcal{D}_{x}}}\right),$ (3)
$\text{PE}_{y}^{(2i)}(y)$ $=\sin\left(\frac{y}{10000^{2i/\mathcal{D}_{y}}}\right),\quad\text{PE}_{y}^{(2i+1)}(y)=\cos\left(\frac{y}{10000^{2i/\mathcal{D}_{y}}}\right),$
$\text{PE}_{z}^{(2i)}(z)$ $=\sin\left(\frac{z}{10000^{2i/\mathcal{D}_{z}}}\right),\quad\text{PE}_{z}^{(2i+1)}(z)=\cos\left(\frac{z}{10000^{2i/\mathcal{D}_{z}}}\right),$
where $i=0,1,\ldots,\mathcal{D}_{x}/2-1$.
@@
$\mathbf{PE}_{\text{encoder}}$ $=\text{stack}\Big(\text{PE}_{\text{[CLS]}},$ (4)
$\quad\left\{\text{concat}\big(\text{PE}_{x}(x),\,\text{PE}_{y}(y),\,\text{PE}_{z}(z)\big)\;\middle|\;(x,y,z)\in\{0,\ldots,\mathcal{G}-1\}^{3}\right\}\Big)\in\mathbb{R}^{(1+N_{p})\times\mathcal{D}_{e}}.$
@@ 2.3.1 Downstream task I: Homogenized stiffness prediction
2.3.2 Formulation of the ODMN
Following the motivation introduced in Section 1, this section summarizes the formulation of the orientation-aware interaction-based deep material network (ODMN) relevant to the present study. The ODMN serves as a physically interpretable surrogate model for polycrystalline materials and constitutes an extension of the deep material network (DMN) architecture [liu2019deep, noels2022micromechanics, noels2022interaction], with explicit incorporation of crystallographic texture effects [Wei01].
Conceptually, the ODMN represents a polycrystalline microstructure as a parameterized reduced-order system that approximates the homogenized mechanical response of a full-field representative volume element (RVE) [wu2025stochastic, shin2023deep]. Starting from a unit cell $\Omega$, the reduced-order microstructure is hierarchically partitioned into $2^{N}$ subdomains,
@@ 2.3.2 Formulation of the ODMN
where each subdomain $\Omega_{i}$ corresponds to a material node $\mathcal{M}^{i}$ in the network. The network depth $N$ controls the resolution of the reduced-order representation, with larger $N$ enabling a more expressive approximation of the homogenized response.
The hierarchical partitioning is encoded by a binary-tree material network of depth $N$. At each internal tree node located at level $l$ and position $p$, an interaction direction $\vec{\mathbf{N}}^{\,l}_{p}$ is introduced to characterize the interface orientation governing stress equilibrium between adjacent subdomains. At each leaf-level material node $\mathcal{M}^{i}$, a set of parameters $\{z^{i},\alpha^{i},\beta^{i},\gamma^{i}\}$ is assigned, representing the volume fraction and local crystallographic orientation of subdomain $\Omega_{i}$.
The interaction direction $\vec{\mathbf{N}}$ is parameterized by two angular variables, $\theta$ and $\phi$, defining a unit normal vector according to
@@
For an ODMN of depth $N$, the complete set of trainable parameters is given by
@@
where the first set corresponds to material-node parameters and the second set specifies the interaction directions at each internal tree node.
A key property of the ODMN is that its parameters can be identified during an offline training stage using only linear elastic homogenized stiffness data, while the resulting network remains capable of extrapolating to nonlinear mechanical behavior during online prediction [liu2019deep, gajek2020micromechanics, noels2022micromechanics, wan2024decoding]. This capability arises because the learned parameters encode a first-order approximation of the microstructural mechanical response under linear elasticity, which is preserved within the hierarchical homogenization operators during online evaluation [gajek2020micromechanics].
During online prediction, the ODMN performs a downscaling operation from the macroscopic deformation gradient $\bar{\mathbf{F}}$ to the deformation gradient $\mathbf{F}^{i}$ of each subdomain $\Omega_{i}$ according to
@@
where the interaction coefficients $\alpha^{i,j}$ and interaction directions $\mathbf{N}^{j}$ are determined by the parameter set $\mathcal{F}$.
At each material node, the local constitutive response is evaluated through
@@
where $\mathbf{P}^{i}$ denotes the first Piola–Kirchhoff stress of the subdomain $\Omega_{i}$, $\partial\mathbf{P}^{i}/\partial\mathbf{F}^{i}$ is the corresponding consistent tangent operator, and $\mathbb{P}^{i}(\cdot)$ represents the local material law parameterized by the internal state variables $\boldsymbol{\psi}^{i}$ (e.g., history-dependent variables in crystal plasticity). The resulting local stresses and tangents are subsequently upscaled through the hierarchical network to recover the homogenized stress $\bar{\mathbf{P}}$ and its consistent tangent.
The interaction vectors $\mathbf{a}^{j}$ represent fluctuation fields associated with the interfaces between adjacent subdomains. During each Newton–Raphson iteration, the interaction vectors are initialized as $\mathbf{a}^{j}=\mathbf{0}$ and updated to satisfy the Hill–Mandel condition
@@
where $W^{i}$ denotes the volume fraction of subdomain $\Omega_{i}$.
Similar to the original DMN framework, ODMN parameters are tied to specific microstructural realizations, which limits direct transferability across different microstructures. In the present work, this limitation is addressed by coupling the ODMN with a foundation model that provides transferable latent representations of polycrystalline microstructures. The resulting framework enables systematic inference of ODMN parameters for previously unseen three-dimensional microstructures, as described in the following subsection.